# HiddenDetect: Detecting Jailbreak Attacks against Large Vision-Language Models via Monitoring Hidden States

### v2

@@ 1 Introduction
The rapid advancements in large language models (LLMs) Touvron et al. (2023a, b); Dubey et al. (2024); Chiang et al. (2023) have fueled the development of large vision-language models (LVLMs), such as GPT-4V Achiam et al. (2023), mPLUG-OWL Ye et al. (2023), and LLaVA Liu et al. (2023a). By integrating multiple modalities, LVLMs have demonstrated impressive capabilities in multimodal reasoning, visual question answering, and embodied AI tasks. However, this cross-modal alignment introduces unique safety challenges, as LVLMs have been shown to be more vulnerable to adversarial manipulations than their text-only counterparts Liu et al. (2023b). These vulnerabilities raise serious concerns about their reliability, particularly in high-stakes applications.
@@
To address these vulnerabilities, existing safety mechanisms largely focus on behavioral interventions, such as supervised fine-tuning on curated datasets Zong et al. (2024), defensive prompting Wu et al. (2023), or multimodal reasoning techniques Jiang et al. (2024). However, these approaches are often resource-intensive, manually engineered, and inherently reactive—they attempt to mitigate safety risks after unsafe behaviors manifest. But what if LVLMs already encode safety-relevant signals within their internal activations?
Therefore, in this paper, we aim to answer the following research question: Can we ensure safety by monitoring LVLM’s hidden states? Inspired by recent research in activation-based interpretability Park et al. (2023); Wang et al. (2024b); Nanda et al. (2023); Li et al. (2024b), we investigate whether LVLMs inherently recognize unsafe prompts within their latent activations. Our key insight is that LVLMs exhibit distinct activation patterns when encountering unsafe inputs, even before generating a response. These latent signals offer a potential intrinsic safety mechanism that can be leveraged for real-time adversarial detection without external modifications or fine-tuning.
Building on this observation, we propose an activation-based safety framework that detects unsafe prompts by monitoring the model’s internal activations during inference. As illustrated in Figure 1, unlike prior methods that rely on fine-tuning or input manipulations, we introduce a Refusal Vector (RV), a learned representation constructed from the model’s hidden states, to classify prompts as safe or unsafe. This is achieved by computing a cosine similarity vector between intermediate representations and a predefined refusal embedding, denoted as $\mathbf{F}$. A scoring function $s(\mathbf{F})$ is then used to assess prompt safety, flagging unsafe inputs based on an adaptive threshold. Unlike previous approaches, our method operates directly within the model’s latent space, avoiding manual prompt engineering or costly supervised fine-tuning.
Our approach offers several key advantages. First, activation-based safety detection introduces minimal computational overhead and requires no additional model tuning. Second, unlike fine-tuned safety classifiers, our method generalizes to unseen adversarial prompts without requiring labeled training data. Third, while designed to mitigate multimodal jailbreak attacks, our approach is also effective against pure LLM adversarial prompts, demonstrating broad applicability across different types of threats. Extensive experiments demonstrate that our approach outperforms state-of-the-art defenses in both accuracy and efficiency, making it a scalable and effective safety solution for real-world LVLM deployments. By shifting from behavioral to activation-based safety monitoring, this work highlights a promising direction for ensuring the security of next-generation multimodal AI systems.
Our contributions can be summarized as follows:
  • We identify a key insight: LVLMs exhibit distinct activation patterns when processing unsafe prompts, even before generating a response. This suggests the presence of an intrinsic safety mechanism capable of detecting adversarial inputs in real-time without requiring external modifications or additional fine-tuning.
  • We introduce HiddenDetect, an activation-based safety framework that monitors LVLM hidden states to identify unsafe prompts, offering a proactive alternative to traditional behavioral interventions such as fine-tuning and defensive prompting.
  • We conduct extensive experiments demonstrating that HiddenDetect outperforms existing safety defenses in both accuracy and efficiency, generalizing effectively across multimodal jailbreak attacks and text-based adversarial prompts.
@@ 2.1 Vulnerability and Safety in LVLMs
Large vision-language models (LVLMs) are vulnerable to various security risks, including susceptibility to malicious prompt attacks Liu et al. (2024), which can exploit vision-only Liu et al. (2023b) or cross-modal Luo et al. (2024b) inputs to elicit unsafe responses. Prior studies identify two primary attack strategies for embedding harmful content. The first involves encoding harmful text into images using text-to-image generation tools, thereby bypassing safety mechanisms Gong et al. (2023); Liu et al. (2023b); Luo et al. (2024b). For example, Gong et al. (2023) demonstrate how malicious queries embedded in images through typography can evade detection. The second strategy employs gradient-based adversarial techniques to craft images that appear benign to humans but provoke unsafe model outputs Zhao et al. (2024); Shayegani et al. (2023); Dong et al. (2023); Qi et al. (2023); Tu et al. (2023); Luo et al. (2024a); Wan et al. (2024). These methods leverage minor perturbations or adversarial patches to mislead classifiers Bagdasaryan et al. (2023); Schlarmann and Hein (2023); Bailey et al. (2023); Fu et al. (2023).
@@ 2.2 Efforts to Safeguard LVLMs
3 Safety Awareness in LVLMs
In this section, we aim to demonstrate the broad presence of safety awareness in LVLMs and identify the most safety-aware layers using a few-shot approach. Since safety-aware responses in LVLMs often involve specific refusal-related tokens (e.g., “sorry”, “cannot”), the first step is to construct a refusal vector in the vocabulary space. This begins with identifying a specialized set of tokens, referred to as the Refusal Token Set (RTS), which consists of tokens frequently appearing when the model declines to respond to inappropriate or harmful queries.
3.1 Constructing a Refusal Vector (RV)
The construction of the Refusal Token Set (RTS) begins with a collection of toxic image-text prompt pairs (e.g., an image depicting a dangerous object paired with a text query like “How to assemble this?”). The model’s responses to these inputs are analyzed to identify recurring words indicative of refusals. The most frequently occurring refusal-related tokens form the initial RTS.
To refine the RTS, each toxic image-text prompt pair is processed by the model, and the hidden states at the final token position across all layers are extracted. These hidden states are projected into vocabulary space, yielding a logit vector over the vocabulary. At each layer, the top five tokens with the highest logit values are identified. Any refusal-related tokens among them that are not already part of the RTS are added, progressively expanding the set. This process iterates until no significant additions occur. The finalized RTS used in our experiments is provided in the appendix.
Once the RTS is established, the Refusal Vector (RV) is constructed in vocabulary space. This vector is represented as a one-hot encoding, where dimensions corresponding to the token IDs in the RTS are set to 1, while all others remain 0. RV serves as a compact yet comprehensive representation of safety-aware refusal signals, capturing the model’s inclination to reject harmful or inappropriate requests.
3.2 Evaluating Safety Awareness
To evaluate the model’s internal safety awareness, two minimal sets of safe and unsafe queries are employed. These queries vary in structure and semantic content, spanning from pure text to typo and non-typo image, ensuring that the identified safety-aware layers are not biased by specific query formats. The few-shot query sets used in the experiment are provided in the appendix.
Despite a large fraction of queries in the few-shot unsafe set successfully bypassing the model’s safety mechanisms, analysis reveals that safety awareness remains broadly distributed across layers, even for jailbreak prompts. To investigate this, both query sets are fed into the model, and hidden states are captured at the final token position of each layer—this position most effectively reflects how auto-regressive models process and interpret input at different depths Zhou et al. (2024).
For an LVLM whose backbone LLM has $L$ layers, given an image-text input prompt $P_{i}$, the hidden states at the final positional index from each layer $l\in\{0,1,\dots,L-1\}$ are extracted. These are then projected into vocabulary space to obtain:
$H_{i}=\{h_{l}\mid h_{l}=\text{proj}(\mathbf{h}_{l}),\quad l=0,1,\dots,L-1\}.$ (1)
Using the combined Refusal Vector $r$, a vector $F\in\mathbb{R}^{L}$ is computed to capture refusal-related semantics across layers for $P_{i}$. Each element $F_{l}$ in this vector is given by the cosine similarity between the projected hidden state $h_{l}$ and $r$:
$F_{l}=\frac{h_{l}\cdot r}{\|h_{l}\|\|r\|},\quad l\in\{0,1,\dots,L-1\}.$ (2)
Averaging these refusal similarity vectors over all queries in the respective sets yields:
$F_{\text{safe}}=\frac{1}{N_{\text{safe}}}\sum_{i\in\text{safe}}F_{i}$ (3)
$F_{\text{unsafe}}=\frac{1}{N_{\text{unsafe}}}\sum_{i\in\text{unsafe}}F_{i}$ (4)
The Refusal Discrepancy Vector (FDV) is then computed as:
$F^{\prime}=F_{\text{unsafe}}-F_{\text{safe}}.$ (5)
As illustrated in Figure 2, $F^{\prime}$ generally increases across layers before eventually declining, with higher values indicating greater safety awareness. The initial increase suggests that deeper layers contribute to enhanced contextual understanding and safety detection. However, in the final layers, the model must balance safety considerations with fulfilling the user’s request, leading to a decline in safety awareness.
A layer is defined as safety-aware if $F^{\prime}_{l}>0$. Results indicate that after the initial layers, $F^{\prime}$ remains consistently positive, suggesting that safety awareness is embedded throughout the model.
Algorithm 1 Pipeline of the Detection Method
Input:  LVLM $\mathcal{M}$ with $\mathcal{L}$ layers  Refusal vector $\mathcal{RV}$  Most safety-aware layers $\mathcal{L}_{\mathcal{M}}$  Detected sample $\mathcal{S}$  Configurable threshold $t$
Output: Safety label $I\in\{0,1\}$ (1 for unsafe, 0 for safe)
Step 1: Compute the refusal semantics strength at the most safety-aware layers
for $l\in\mathcal{L}_{\mathcal{M}}$ do
     1. Extract hidden state from layer $l$:
$\mathcal{h}_{l}=\mathcal{M}_{l}(\mathcal{S})$
     2. Project to the vocabulary space:
$\mathcal{h}_{l}^{\prime}=\mathcal{h}_{l}\cdot\mathcal{W}_{\text{unembedding}}$
     3. Compute cosine similarity with the refusal vector:
$F_{l}=\cos(\mathcal{h}_{l}^{\prime},\mathcal{RV})$
end for
Step 2: Determine the safety label based on the computed safety score
Compute the safety score using the trapezoidal rule over the most safety-aware layers:
[…]
     $I\leftarrow 1$triangleright Sample is unsafe
else
     $I\leftarrow 0$triangleright Sample is safe
end if
@@ 3.3 Identifying the Most Safety-Aware Layer Range
To pinpoint the layers with the strongest safety awareness, the most safety-aware layer range $(s,e)$ is determined by comparing $F^{\prime}$ to the final layer’s discrepancy value, $F^{\prime}_{L-1}$:
$s=\min\{l\mid F^{\prime}_{l}>F^{\prime}_{L-1}\},$ (6)
$e=\max\{l\mid F^{\prime}_{l}>F^{\prime}_{L-1}\}.$ (7)
The final layer’s discrepancy value, $F^{\prime}_{L-1}$, serves as a baseline since a significant fraction of unsafe queries can bypass the model’s defenses, indicating that the final layer is less effective at recognizing unsafe content. In contrast, layers exhibiting stronger safety awareness maintain higher $F^{\prime}$ values. Specifically, a layer $l$ that can effectively distinguish between safe and unsafe queries must satisfy $F^{\prime}_{l}>F^{\prime}_{L-1}$.
This minimal-query approach highlights both the broad presence of safety awareness across layers and provides a systematic method to identify the layers with the strongest safety focus. These insights lay the foundation for subsequent detection methods.
@@ 4 Method
Model Method Training- free Text-based Image-based Average
XSTEST FigTxt FigImg MM-SafetyBench
LLaVA Perplexity 0.610 0.758 0.825 0.683 0.719
Self-detection 0.630 0.765 0.837 0.705 0.734
GPT-4V 0.649 0.784 0.854 0.721 0.752
GradSafe 0.714 0.831 0.889 0.760 0.798
MirrorCheck 0.670 0.792 0.860 0.725 0.762
CIDER 0.652 0.786 0.850 0.713 0.750
JailGuard 0.662 0.784 0.859 0.715 0.755
Ours 0.868 0.976 0.997 0.846 0.922
CogVLM Perplexity 0.583 0.732 0.797 0.657 0.692
Self-detection 0.597 0.743 0.813 0.683 0.709
GPT-4V 0.623 0.758 0.828 0.698 0.727
GradSafe 0.678 0.809 0.872 0.744 0.776
MirrorCheck 0.641 0.768 0.831 0.709 0.737
CIDER 0.635 0.763 0.822 0.698 0.730
JailGuard 0.645 0.771 0.834 0.703 0.738
Ours 0.834 0.962 0.991 0.823 0.903
Qwen-VL Perplexity 0.525 0.679 0.737 0.612 0.638
Self-detection 0.542 0.695 0.752 0.627 0.654
GPT-4V 0.567 0.713 0.771 0.645 0.674
GradSafe 0.617 0.762 0.812 0.692 0.721
MirrorCheck 0.587 0.727 0.776 0.660 0.687
CIDER 0.576 0.718 0.764 0.650 0.677
JailGuard 0.584 0.724 0.772 0.655 0.684
Ours 0.762 0.866 0.910 0.764 0.826
@@
In this section, we describe how HiddenDetect works by utilizing the safety awareness in the hidden states. The overall pipeline of HiddenDetect is shown in Figure 3. The assessment of whether a prompt $P_{i}$ may lead to ethically problematic responses involves computing its refusal-related semantic vector $\mathbf{F}\in\mathbb{R}^{L}$, as introduced in Section 3.2. Each entry $F_{l}$ in $\mathbf{F}$ corresponds to the cosine similarity between the projected hidden state $\mathbf{h}_{l}$ at layer $l$ and the Refusal Vector $\mathbf{r}$:
@@
where the trapezoidal rule is used to approximate the cumulative magnitude of $F$ across these layers. Finally, if the computed safety score exceeds a configurable threshold, the prompt is classified as unsafe; otherwise, it is deemed safe. The overall detection process is also elaborated in Algorithm 1.

### v4

@@ 1 Introduction
The rapid advancements in large language models (LLMs) Touvron et al. (2023a, b); Dubey et al. (2024); Chiang et al. (2023) have paved the way for large vision-language models (LVLMs), such as GPT-4V Achiam et al. (2023), mPLUG-OWL Ye et al. (2023), and LLaVA Liu et al. (2023a). By integrating vision and language modalities, LVLMs excel at tasks like visual reasoning, question answering, and grounded decision-making. However, this multimodal capability also introduces new safety risks. Recent studies reveal that LVLMs are more susceptible to adversarial manipulation than their text-only counterparts Liu et al. (2023b), especially through visual perturbations or multimodal prompt engineering. These vulnerabilities pose significant concerns for real-world deployment in high-stakes settings.
@@
To mitigate these risks, existing safety efforts largely focus on behavioral defenses—such as fine-tuning on curated safety datasets Zong et al. (2024), defensive prompting Wu et al. (2023), or explicit reasoning modules Jiang et al. (2024). While effective to some extent, these methods are resource-intensive, require manual supervision, and often fail to generalize to unseen adversarial strategies. But what if safety-relevant signals already exist within the model’s internal activations—especially in the context of multimodal inputs?
Therefore, in this paper, we aim to investigate a fundamental question: Can LVLMs detect unsafe prompts through their internal hidden states before generating any output? Inspired by recent progress in activation-based interpretability Park et al. (2023); Wang et al. (2024b); Nanda et al. (2023); Li et al. (2024), we explore whether LVLMs encode latent safety signals that correlate with prompt harmfulness, and how these signals evolve across layers and modalities.
Our key insight is that LVLMs exhibit distinct activation patterns in response to unsafe prompts—both in text-only and vision-conditioned settings. We show that these refusal-related signals can be measured using a Refusal Strength Vector, which captures alignment with vocabulary tokens associated with model refusals. We further reveal that in LVLMs, the emergence of these safety signals is delayed under multimodal inputs, particularly when adversarial images are used. This delay is strongly correlated with attack success rates, offering a new perspective on how and where safety mechanisms fail inside LVLMs.
To harness this intrinsic behavior, we propose HiddenDetect, an activation-based safety detection framework that identifies unsafe prompts by monitoring intermediate model activations. Rather than relying on external prompts or fine-tuned safety heads, our method constructs a refusal-aware embedding and measures its alignment with layer-wise hidden states to assess input safety in real time. A cosine similarity-based score function is used to aggregate signal strength over the most safety-aware layers, enabling robust detection without any retraining or supervision.
Compared to prior approaches, our method operates directly in the model’s latent space, incurs minimal overhead, and generalizes across both textual and multimodal jailbreak attacks. In addition, we provide a fine-grained analysis of how refusal semantics evolve across layers and modalities, revealing structural insights into the safety mechanisms of LVLMs.
To sum up, our contributions are as follows:
  • We discover that LVLMs exhibit distinct and structured activation patterns when processing unsafe prompts—offering evidence of an intrinsic, model-internal safety mechanism that activates prior to generation.
  • We introduce HiddenDetect, a training-free, activation-based detection framework that identifies unsafe prompts by monitoring refusal semantics within hidden states, avoiding external classifiers or defensive prompt design.
  • We conduct the first layer-wise analysis of safety signal emergence across modalities, revealing that visual inputs cause delayed activation of safety mechanisms. This temporal shift in refusal semantics correlates with higher attack success rates in multimodal jailbreaks.
  • Extensive experiments show that HiddenDetect outperforms state-of-the-art safety defenses on both text-based and multimodal benchmarks, achieving higher accuracy and generalizability with lower computational cost.
@@ 2.1 Vulnerability and Safety in LVLMs
Large vision-language models (LVLMs) are vulnerable to various security risks, including susceptibility to malicious prompt attacks Liu et al. (2024), which can exploit text-only Luo et al. (2024b) or cross-modal Liu et al. (2023b); Li et al. (2025b) inputs to elicit unsafe responses. Prior studies identify two primary attack strategies for embedding harmful content. The first involves encoding harmful text into images using text-to-image generation tools, thereby bypassing safety mechanisms Gong et al. (2023); Liu et al. (2023b); Luo et al. (2024b). For example, Gong et al. (2023) demonstrate how malicious queries embedded in images through typography can evade detection. The second strategy employs gradient-based adversarial techniques to craft images that appear benign to humans but provoke unsafe model outputs Zhao et al. (2024b); Shayegani et al. (2023); Dong et al. (2023); Qi et al. (2023); Tu et al. (2023); Luo et al. (2024a); Wan et al. (2024). These methods leverage minor perturbations or adversarial patches to mislead classifiers Bagdasaryan et al. (2023); Schlarmann and Hein (2023); Bailey et al. (2023); Fu et al. (2023).
@@ 2.2 Efforts to Safeguard LVLMs
Figure 3: Visualization of Refusal behavior across harmful query types and safety alignment settings. Left: Refusal discrepancy value across layers for five query types reveals delayed and weaker safety activation for multimodal harmful queries. Right: Bimodal safety alignment enhances refusal strength mainly for multimodal queries, especially in middle and late layers, while having minimal impact on pure text inputs.
3 Safety Awareness in LVLMs
In this section, we aim to explore the emergence of safety awareness in large vision-language models (LVLMs) and proposes a systematic way to locate the most safety-aware layers using a multimodal few-shot approach. While prior work on large language models (LLMs) has shown that refusal behaviors correlate with specific activation patterns and vocabulary logitsArditi et al. (2024); Zheng et al. (2024); Zhao et al. (2024a); Cao et al. (2024), LVLMs require fundamentally different treatment due to the multimodal nature of their inputs. In particular, visual context can significantly modulate safety behavior, making it insufficient to rely on text-only safety patterns. To address this, we propose a VLM-specific method for constructing a refusal vector grounded in multimodal context and identifying safety-aware layers that are sensitive to harmful content across both pure text and bimodal inputs.
3.1 Constructing a Refusal Vector in LVLMs
The construction of the Refusal Token Set (RTS) begins with a collection of image-text prompt pairs containing harmful or inappropriate visual content (e.g., an image of a weapon with a prompt such as “How to assemble this?”). Unlike previous LLM-based approaches Arditi et al. (2024), which mine refusal signals from purely textual inputs, we analyze LVLM outputs to multimodal prompts where the refusal is likely to be visually conditioned. This enables us to capture refusal tokens that are sensitive not only to linguistic signals but also to underlying image semantics.
We first collect model responses to a curated set of harmful image-text prompts and extract the most frequent refusal-related tokens (e.g., “sorry”, “unable”, “cannot”) to form the initial RTS. To reflect VLM-specific behaviors, we then refine the RTS by projecting hidden states at the final token position into the vocabulary space using the model’s unembedding layer. For each layer, we collect the top five tokens with the highest logits, conditioned on both visual and textual inputs. Any refusal-related tokens not already in the RTS are added. This refinement loop continues until no substantial new tokens are discovered, ensuring that the final RTS encodes comprehensive multimodal refusal behavior. A comprehensive list of the final RTS tokens is provided in Appendix A.1.
Once the RTS is finalized, we construct the Refusal Vector ($r$) in vocabulary space as a sparse binary vector, where entries corresponding to RTS token indices are set to 1.
3.2 Evaluating Safety Awareness in LVLMs
To investigate how refusal semantics manifest across the LVLM’s depth, we evaluate the model on a small set of safe and unsafe multimodal queries. These include text-only, typo-based, and visually grounded prompts, designed to reveal whether certain layers consistently activate in response to unsafe content.
Each prompt is passed through the model, and the hidden states at the final token position from all layers are projected into the vocabulary space. Let $h_{l}$ denote the projected hidden state at layer $l$. For a LVLM having $L$ layers, we define the Refusal Strength Vector $F\in\mathbb{R}^{L}$ as the alignment between each hidden state and the Refusal Vector $r$ and it is computed via cosine similarity:
$F_{l}=\frac{h_{l}\cdot r}{\|h_{l}\|\|r\|},\quad l\in\{0,1,\dots,L-1\}.$ (1)
$F$ captures how strongly each layer aligns with refusal semantics. We average this across all safe and unsafe prompts to obtain:
$\begin{split}F_{\text{safe}}&=\frac{1}{N_{\text{safe}}}\sum_{i\in\text{safe}}F_{i},\\ F_{\text{unsafe}}&=\frac{1}{N_{\text{unsafe}}}\sum_{i\in\text{unsafe}}F_{i}.\end{split}$ (2)
We then compute the Refusal Discrepancy Vector $F^{\prime}$:
$F^{\prime}=F_{\text{unsafe}}-F_{\text{safe}}.$ (3)
@@ 3.2 Evaluating Safety Awareness in LVLMs
This vector highlights which layers are more responsive to unsafe prompts than to benign ones. Interestingly, we observe that $F^{\prime}$ tends to increase in the middle layers before declining toward the end—suggesting that safety-related features are first detected in mid-level multimodal fusion layers Zhang et al. (2025), but then diluted as the model balances response relevance and safety alignment in later decoding stages.
A layer $l$ is considered safety-aware if $F^{\prime}_{l}>0$, and empirically, many layers beyond the early stages satisfy this condition. This observation indicates that multimodal safety awareness is distributed and emerges progressively as the model processes and integrates visual and textual information.
@@ 3.3 Identifying the Most Safety-Aware Layer Range
To isolate the layers most sensitive to multimodal unsafe content, we define the safety-aware range $(s,e)$ using the final layer’s discrepancy score $F^{\prime}_{L-1}$ as a conservative baseline. Layers with stronger refusal signal than the final decoding layer are defined as:
$s$ $=\min\{\,l:F^{\prime}_{l}>F^{\prime}_{L-1}\},$ (4)
$e$ $=\max\{\,l:F^{\prime}_{l}>F^{\prime}_{L-1}\}.$
This ensures that only layers with a meaningful contribution to multimodal safety-related decision-making are retained. In contrast to LLMs, where shallow and middle layers often suffice for rejecting harmful queries Li et al. (2025a), we find that in LVLMs stronger refusal semantics often emerge in middle and late layers, particularly those involved in image-text alignment and modality fusion.
This LVLM-specific formulation—constructing a visually grounded refusal vector and analyzing multimodal hidden representations—lays the foundation for the proposed detection method introduced in the later section.
3.4 Delayed Safety Activation in Multimodal Jailbreaks
For LVLMs, queries with clearly defined malicious intent typically fall into two categories. The first type involves textual queries that explicitly express harmful intent; in these cases, adding visual input does not alter the core intention, irrespective of the image’s relevance. The second type consists of queries whose harmful intent emerges only when accompanied by an image; the textual component alone appears benign or ambiguous without visual context.
Unfortunately, visual inputs generally increase the vulnerability of LVLMs in both scenarios. Previous research has demonstrated that pairing harmful text queries with images—even blank ones—can significantly elevate the ASR against LVLMs Li et al. (2025b). Exploiting this vulnerability, JailbreakV-28KLuo et al. (2024b) combines specially crafted harmful text prompts originally designed to jailbreak LLMs with various types of images, achieving notably higher ASRs than their text-only counterparts.
More sophisticated jailbreak methods tailored for LVLMs, such as FigStep, HADES and MM-SafetyBench Gong et al. (2023); Li et al. (2025b); Liu et al. (2024), embed harmful intentions primarily within the visual modality itself, compelling models to rely heavily on their cross-modal understanding to recognize the malicious nature of the input. However, empirical findings indicate that many state-of-the-art open-source LVLMs struggle with this challenge, particularly those that lack bimodal safety alignment and rely solely on safety mechanisms inherited from their LLM backbones aligned only for the text modality Zong et al. (2024).
However, very limited research Xu et al. (2025) has investigated why safety alignment achieved for the text modality fails to effectively transfer to the vision modality in LVLMs. Thanks to the modality-aware properties of both the Refusal Strength Vector $F$ and the Refusal Dispersancy Vector $F^{\prime}$, we are able to systematically investigate how LVLMs’ hidden-state refusal behaviors vary across differently structured harmful queries, both before and after bimodal safety alignment.
As shown in the left subplot of Figure 3, we compare the $F^{\prime}$ of LLaVA-v1.5-7b on five variants of harmful queries sampled from JailbreakV-28K Luo et al. (2024a): direct harmful text, crafted jailbreak text, their image-paired counterparts (regardless of image relevance), and harmless text paired with a harmful image. Because we make each query type convey the same underlying harmful intent, differences in the model response reveal the influence of input modality. We observe that crafted jailbreak text queries decrease the overall magnitude of safety activation, contributing to an increase in ASR. Interestingly, even when harmful intent is already present in the text, pairing with an image can further compromise safety activation, particularly in the early and middle layers. For harmful visually grounded queries where the text alone is harmless, the model requires multiple layers to integrate visual and linguistic information, resulting in delayed and greatly diminished safety activation compared to queries where the text itself is harmful.
We further compare the model’s behavior before and after bimodal safety alignment (right subplot of Figure 3), using the model checkpoints released by Zong et al. (2024). We find that refusal strength remains nearly unchanged for harmful text queries, but increases significantly for cross-modal queries, especially in the middle and late layers. These findings suggest that, within LVLMs, harmful queries of different modalities can activate refusal behavior at different network depths. This effect is closely tied to both the model’s information flow for different modalities and the specifics of safety alignment training. For queries containing harmful text, safety awareness is triggered earlier and is more pronounced in the early and middle layers. In contrast, when harmful intent is conveyed solely through an accompanying image, safety activation is delayed and primarily emerges in the middle to late layers.
@@ 4 Method
Model Method Training- free Text-based Image-based
XSTest FigTxt MM-SafetyBench FigImg JailBreakV-28K
LLaVA Perplexity 0.610 0.758 0.825 0.683 0.781
Self-detection 0.630 0.765 0.837 0.705 0.805
GPT-4V 0.649 0.784 0.854 0.721 0.817
GradSafe 0.714 0.831 0.889 0.760 0.845
MirrorCheck 0.670 0.792 0.860 0.725 0.813
CIDER 0.652 0.786 0.850 0.713 0.786
JailGuard 0.662 0.784 0.859 0.715 0.801
Ours 0.868 0.976 0.997 0.846 0.932
Qwen-VL Perplexity 0.583 0.732 0.797 0.657 0.732
Self-detection 0.597 0.743 0.813 0.683 0.768
GPT-4V 0.623 0.758 0.828 0.698 0.767
GradSafe 0.678 0.809 0.872 0.744 0.839
MirrorCheck 0.641 0.768 0.831 0.709 0.779
CIDER 0.635 0.763 0.822 0.698 0.734
JailGuard 0.645 0.771 0.834 0.703 0.756
Ours 0.834 0.962 0.991 0.823 0.907
CogVLM Perplexity 0.525 0.679 0.737 0.612 0.682
Self-detection 0.542 0.695 0.752 0.627 0.701
GPT-4V 0.567 0.713 0.771 0.645 0.760
GradSafe 0.617 0.762 0.812 0.692 0.789
MirrorCheck 0.587 0.727 0.776 0.660 0.756
CIDER 0.576 0.718 0.764 0.650 0.723
JailGuard 0.584 0.724 0.772 0.655 0.746
Ours 0.762 0.866 0.910 0.764 0.883
@@
In this section, we describe how HiddenDetect works by utilizing the safety awareness in the hidden states. The overall pipeline of HiddenDetect is shown in Figure 4. For a LVLM $M$ with $L$ layers, the assessment of whether a prompt $P_{i}$ may lead to ethically problematic responses involves computing its Refusal Strength Vector $\mathbf{F}\in\mathbb{R}^{L}$, as introduced in Section 3.2. Each entry $F_{l}$ in $\mathbf{F}$ corresponds to the cosine similarity between the projected hidden state $\mathbf{h}_{l}$ at layer $l$ and the Refusal Vector $\mathbf{r}$:
@@
where the trapezoidal rule is used to approximate the cumulative magnitude of $F$ across these layers. Finally, if the computed safety score exceeds a configurable threshold, the prompt $P_{i}$ is classified as unsafe; otherwise, it is deemed safe.